OpenStudy (anonymous):

Solve using two different methods. Explain which method you found to be more efficient. A. 3x-9y=3 6x-3y=-24 B. 7x-3y=20 5x+3y=16 C. y=1/2x-6 2x+6y=19 Please help me and show your work!! Thank you! @agent0smith @radar @campbell_st @helpme1.2 @bibby @mebs

OpenStudy (usukidoll):

there's elimination method, subsitution, and for the really advanced yet super eay, matrix method

OpenStudy (anonymous):

Okay, can we do elimination and matrix?

OpenStudy (usukidoll):

sure.. I wish draw would still work on here.. twiddla is the alternative

OpenStudy (anonymous):

Okay..So how do I do it?

OpenStudy (usukidoll):

http://www.twiddla.com/1470390 draw the problems ... equations from here

OpenStudy (anonymous):

We are going to work through twiddla?

OpenStudy (anonymous):

@agent0smith can you help me?

OpenStudy (usukidoll):

sorry I was kind of busy with my own probz... but I did solve the first one through matrix on twiddla

OpenStudy (anonymous):

I don't see it

OpenStudy (usukidoll):

http://www.twiddla.com/1470390

OpenStudy (usukidoll):

zoom out

OpenStudy (usukidoll):

control - to zoom out

OpenStudy (anonymous):

I don't understand that! I am sorry, but it looks like a bunch of gibberish. Can you make some kind of chart here?

OpenStudy (usukidoll):

os isn't working too well for me at the moment

OpenStudy (anonymous):

I need someone who can help me do these problems, @UsukiDoll is there anyway that you can make a chart on anything to help me see it?

OpenStudy (usukidoll):

I don't think I could if os is messing with me at the moment. :(

OpenStudy (anonymous):

Can anyone help me?

OpenStudy (texaschic101):

SUBSTITUTION : 3x - 9y = 3 (when dividing by 3, reduces to x - 3y = 1) x - 3y = 1 --add 3y to both sides x = 3y + 1 now we can use substitution. Sub in 3y + 1 in for x in the other equation 6x - 3y = -24 6(3y + 1) - 3y = -24 -- distribute through the parenthesis 18y + 6 - 3y = -24 -- subtract 6 from both sides 18y - 3y = -24 - 6 -- combine like terms 15y = - 30 -- divide both sides by 15 y = -30/15 y = -2 now sub -2 in for y in either of the equations to find x 3x - 9y = 3 3x - 9(-2) = 3 3x + 18 = 3 3x = 3 - 18 3x = - 15 x = -15/3 x = -5 solution is (-5,-2) ELIMINATION : 3x - 9y = 3 --->(-2) 6x - 3y = -24 ---------------- -6x + 18y = -6 (result of multiplying by -2) 6x - 3y = -24 ----------------add 15y = -30 -- divide by 15 y = -30/15 y = -2 now sub -2 in for y in either of the equations to find x 6x - 3y = -24 6x - 3(-2) = -24 6x + 6 = -24 6x = -24 - 6 6x = -30 x = -30/6 x = -5 Solution is (-5,-2) ======================== ELIMINATION : 7x - 3y = 20 5x + 3y = 16 --------------add 12x = 36 -- divide by 12 x = 36/12 x = 3 now sub 3 in for x in either of the equations to find y 7x - 3y = 20 7(3) - 3y = 20 21 - 3y = 20 -3y = 20 - 21 -3y = -1 y = 1/3 check.. 5x + 3y = 16 5(3) + 3(1/3) = 16 15 + 1 = 16 16 = 16 (correct) solution is (3,1/3) SUBSTITUTION : 7x - 3y = 20 7x = 3y + 20 x = 3/7y + 20/7 now sub 3/7y + 20/7 in for x in the other equation to find y 5x + 3y = 16 5(3/7y + 20/7) + 3y = 16 15/7y + 100/7 + 3y = 16 -- multiply entire equation by 7 15y + 100 + 21y = 112 -- subtract 100 from both sides 15y + 21y = 112 - 100 36y = 12 -- divide by 36 y = 12/36 y = 1/3 now sub 1/3 in for y in either of the equations to find x 7x - 3y = 20 7x - 3(1/3) = 20 7x - 1 = 20 7x = 20 + 1 7x = 21 x = 21/7 x = 3 Solution is (3, 1/3) ===================== SUBSTITUTION : y = 1/2x - 6 now sub 1/2x - 6 in for y in the other equation 2x + 6y = 19 2x + 6(1/2x - 6) = 19 -- distribute through the parenthesis 2x + 3x - 36 = 19 -- add 36 to both sides 2x + 3x = 19 + 36 -- combine like terms 5x = 55 x = 55/5 x = 11 now sub 11 in for x in either of the equations to find y y = 1/2x - 6 y = 1/2(11) - 6 y = 11/2 - 6 --- common denominator is 2 y = 11/2 - 12/2 y = -1/2 solution is (11,-1/2) ELIMINATION : y = 1/2x - 6 -- subtract 1/2x from both sides -1/2x + y = -6 --- multiply by -1 to make x positive 1/2x - y = 6 now we have : 1/2x - y = 6 --->(-4) 2x + 6y = 19 -------------- -2x + 4y = -24 (result of multiplying by -4) 2x + 6y = 19 -------------add 10y = - 5 y = -5/10 y = -1/2 now sub -1/2 in for y in either of the equations to find x 2x + 6y = 19 2x + 6(-1/2) = 19 2x - 3 = 19 2x = 19 + 3 2x = 22 x = 22/2 x = 11 solution is (11,-1/2)